OpenStudy (anonymous):

Can you solve 2(3 – 2x) = x – 9. I don't really understand so could you show your work please?

jimthompson5910 (jim_thompson5910):

2(3-2x)=x-9 6-4x=x-9 -4x=x-9-6 -4x-x=-9-6 -5x=-9-6 -5x=-15 x=-15/(-5) x=3